# A Survey on Brain-inspired Deep Learning via Predictive Coding

### v2

@@ 3.3 Biased Competition and Divisive Input Modulation (BC-DIM)
Figure 5: On the left is an arbitrary neural circuit (non-linearities are omitted for simplicity) that could be constructed under one of the three PC frameworks surveyed. Rao & Ballard PC (PC) only employs generative synapses $\mathbf{W}^{\ell}$, while NGC further employs separate feedback synapses (which may or may not correspond symmetrically to the generative pathway). Note that NGC/BC-DIM notably often employ skip connections ($\mathbf{W}^{7}$). Finally, the right (which zooms in to a single value-error-value node triplet sub-circuit) depicts core differences in internal operations employed by PC/NGC and BC-DIM: PC/NGC computes error nodes via subtraction and updates value nodes using addition, while BC-DIM computes error nodes via division and updates value nodes using multiplication.
While the formulation of biased competitive PC utilizes synaptic update rules similar to PC/NGC, the divisive modulative form requires rules that respect the new types of error units being used. Specifically, the synaptic update to the neural model above would be:
$\Delta\mathbf{W}^{\ell}$ $=\mathbf{W}^{\ell}\odot\Big{(}\beta(\mathbf{z}^{\ell}\cdot((\mathbf{e}^{\ell-1})^{T}-1))\big{)}$ (20)
where we have further factorized the rule in [80] to make it compatible wíth the formulation of the standard PC synaptic update. The full computational instantiation of PC that leverages the above synaptic update as well as the error unit computation and state activity dynamics in Equations (17)–(19) is called the biased competition divisive input modulation (BC-DIM) form of PC.
In addition to enforcing the bio-physical constraint that both synaptic values and neural activities are strictly positive (by design), BC-DIM rests on several useful theoretical properties given its roots in one form of non-negative matrix factorization [81]. Specifically, the error units of BC-DIM computationally entail a minimization of the Kullback-Leibler (KL) divergence between sensory input targets and local model reconstructions. In addition, we can interpret one layer of BC-DIM probabilistically: we equate the prior ($p(H)$) with its hidden state activity (any neuron in $\mathbf{x}^{\ell}_{t}$), evidence ($p(D)$) with the input activity ($\mathbf{x}^{\ell-1}_{t}$), the posterior ($p(H|D)$) with a single synapse inside of $\mathbf{W}^{\ell}$, and the likelihood ($p(D|H)$) with a corresponding feedback synapses (or a single synaptic value inside $(\mathbf{W}^{\ell})^{T}$). This means that the relationship between input activity (in layer $\ell-1$) and the steady-state hidden node activity (in layer $\ell$) is consistent with Bayes’ theorem [80]. Furthermore, the competition implicit in BC-DIM’s neural activities could be viewed as performing a form of “explaining away” [82]: a node winning the competition within a layer could be viewed as the selection of a hypothesis that can explain the input, while suppressing (or reducing the support for) alternative hypotheses.
@@ 4 Predictive Coding in Machine Learning
So far, we have considered the reasons why PC might play a key role in machine intelligence. In this section, we discuss previous efforts that speak to empirical successes in this direction. Specifically, we review lines of work that have tackled different tasks in machine learning. Some of this work uses algorithms that differ in detail from the formalism of Section 3, but still conform to the definition of PC, in spirit.
@@
The first application of PC to supervised learning involved training a small PC network to perform image classification on the MNIST dataset, achieving test and train errors comparable to those of an MLP of the same complexity (depth and width) [46]. Since then, similar results have been achieved on convolutional networks trained on datasets of RGB images, such as CIFAR10 and SVHN [48]. In this work, the authors showed that updating the weight parameters alongside the neural activities (i.e., running Equations 5 and 6 in parallel), instead of waiting for the inference phase to converge, improves test accuracies as well as ensures a better convergence. These results also extend to structured datasets and graph neural networks, where PC is again able to match the performance of backprop on different benchmarks, with the advantage of learning models that are better calibrated and more robust to adversarial examples [51].
@@
Gaussian assumptions (that underwrite PC) can be limiting in scenarios where we need to model different distributions, such as categorical distributions or mixture models. One such scenario is found in transformer models [83]: the attention mechanism encodes a categorical distribution, computed via the softmax activation. To this end, it is possible to generalize the definition of the energy of a specific layer, originally defined as the squared distance between the predictions $\mathbf{u}^{\ell}$ and the activities $\mathbf{x}^{\ell}$, to the KL divergence between two probability distributions, with prediction and activities serving as sufficient statistics. In detail, the PC model now has the following variational free energy:
@@
where $\mathcal{X}^{\ell}(x)$ is a probability distribution defined for every layer with sufficient statistics $x$. If every distribution is a Gaussian, this is equivalent to the energy defined in Equation 4. This generalization allows predictive-coding-based transformers to perform almost as well as standard transformers with the same model complexity [84]. In computational neuroscience, the use of hybrid generative models (with discrete and continuous state spaces) is fairly established: see, e.g., [85] for a first principles account of the requisite variational message passing, which can be regarded as a generalisation of PC within the setting of well-defined hierarchical generative models with mixed discrete and continuous states.
@@
Figure 6: Various systems based on NGC-based PC: (a) a temporal NGC model (orange diamonds depict error nodes, grey circles depict state nodes, while arrows represent synaptic projections) for processing temporally-varying sequences of data patterns, e.g., frames from the moving MNIST dataset, (b) a convolutional NGC model (figure adapted from [75]; green grid blocks represent state feature maps, blue grid blocks represent prediction maps, yellow-orange grid blocks represent error maps, and blue parallelograms depict deconvolutional pathways) for extracting distributed representations of natural images, and (c) the CogNGen architecture (figure adapted from [86]), a modular cognitive control system that implements several core elements of the Common Model of Cognition [87] with PC and vector-symbolic memory ($\mathbf{r}$ depicts a reward, while $\mathbf{a}$ depicts an action vector).
@@
When processing images, PC has shown promise, particularly, when generalizations that utilize convolutional operators are considered [88, 89]. Notably, BC-DIM has been shown to work well at extracting higher-level features from natural images and patches when the PC scheme employs convolution instead of matrix multiplications [90]. Other promising variants of PC have tackled certain sub-problems in vision, including object recognition and discrimination (using over-complete sparse representations) [91, 92].
With respect to NGC, early work relied on simplifying assumptions, particularly that images were grey-scale (or simply single-channel) and exhibited relatively low scene complexity, e.g., no backgrounds and objects were easily identifiable. Note that NGC/PC circuits can be learned with patches instead, where a natural image is deconstructed into a $P\times P$ pixel grids; this allows the predictive synapses to acquire low-level patterns (such as strokes or edges for handwritten digits). Nevertheless, NGC has been generalized to natural images by directly integrating deconvolution/convolution [75], the standard operations used to build convolutional/deconvolutional neural networks, in a framework known as convolutional neural generative coding (Conv-NGC). In Conv-NGC, forward and feedback synapses are implemented to manipulate feature maps, or small clusters of neurons, which are coupled to equally shaped clusters of error units that compute the relevant mismatch signals needed to conduct iterative inference. Surprisingly, Conv-NGC models were shown to learn feature representations that implicitly embodied an image pyramid [93].
@@
PC has enjoyed some success in the much more difficult realm of temporal sequence modeling. Notably, [94, 89, 95, 96] developed a hierarchical neural generative modeling framework based on PC that focused on processing natural video data, where top-down information (inferred latent causal factors) was used to modulate the activities of lower-layers, aiming to extract locally invariant representations. Preceding this development was PC formulated as a Kalman filter [97, 98], which was demonstrated to work well on small sequential snippets of natural images, and deep hierarchies of local recurrent networks [99, 100] shown to work on a small symbol chunking task.
One of the earliest and first formulations of NGC was with respect to time-varying data, i.e., sequences made up of words/sub-words in a sentence or frames of a video [101]. This type of neural system was called the temporal neural coding network (TNCN) and was originally designed to operate as a top-down recurrent generative model and then later generalized to function as a hybrid recurrent and auto-regressive system [68], i.e., the parallel temporal neural coding network (P-TNCN). TNCN and P-TNCN were both shown to process pattern sequences without unfolding (over time), a core requirement for properly training deep recurrent networks (via backprop through time); thus making it local in space as well as time. Notably, both models (on different video simulations/benchmarks) were shown to generalize to unseen/out-of-distribution pattern sequences, exhibiting zero-shot capability, specifically with respect to video data, e.g., TNCN was shown to adapt to bouncing ball videos of different object quantities, while P-TNCN was shown to be capable of synthesizing frames of bouncing digits or characters of different quantities and qualities (without altering synaptic weights on the new samples). More recently, a formulation of PC, called dynamic PC [102], was developed by combining modeling tools from deep learning (i.e., using backprop-based neural circuitry such as hyper-networks) as well as mixture models. This dynamic PC system was demonstrated to acquire useful bio-physically plausible receptive fields in the context of small controlled sequential visual perception problems and tasks.
It is interesting to note that machine learning implementations of PC have been slow to deal with temporal data. This is surprising because PC can be viewed as an extended Kalman filter that should, in principle, handle temporal data in an efficient and Bayes optimal fashion. The generalization of PC as a Bayesian filter that encompasses learning leads to generalized filtering, in which model parameters can be treated as slowly varying latent variables or states. There are many instances of this form of PC; see [103] for a consistent variational treatment. One reason that machine learning may have yet to fully leverage PC in this context is its focus on classification problems that preclude dynamics. To deploy PC for time series assimilation and learning, it is necessary to move from static generative models to state-space models that have explicit dynamics, of the sort touched upon by [102] and implicit in recurrent neural networks.
@@
Another promising aspect of P-TNCNs described above was its ability/potential to conduct continual sequential learning, where [68] demonstrated that its generative ability on previously seen sequence modeling tasks did not degrade/deteriorate nearly as much as in recurrent networks (such as those based on gated recurrent units and long short-term memory cells). This meant that P-TNCNs exhibited some robustness to the catastrophic forgetting that appeared in recurrent networks when learning sequences of different types of patterns. This included recurrent networks trained with backprop but also with online forward mode training schemes that had access to full Jacobians, such as real-time recurrent learning (RTRL) and its noisy first-order approximation unbiased online recurrent learning. Additional effort [67] explored strengthening NGC’s memory retention ability by examining the challenging problem of online cumulative learning, where datasets (or tasks) were presented to the system in the form of a stream and with no indication of when the task was switched (creating a non-stationary, task boundary free variant of the lifelong machine learning problem setup). Surprisingly, the NGC-based model, without access to any task boundaries in [67] (namely, the sequential neural coding network (S-NCN)) was shown to outperform or be competitive with a wide swath of backprop-based approaches that relied on memory-buffers (replay), regularization, and/or auxiliary generative modeling in order to preserve prior knowledge. Stronger performance was seen when another neural circuit drove the S-NCN’s lateral recurrent synapses based on competitive learning (emulating the information routing behavior of the basal ganglia). This simple task mediating circuit was later improved in follow-up effort [104]). Another later effort developed a stochastic implementation of PC with readable/writable memory (BayesPCN) [58] that was shown to be robust to sample-level forgetting.
Active Inference and Robotics.
In computational neuroscience, early formulations of active inference were based on equipping PC with reflexes in order to simulate a variety of behaviors; ranging from handwriting and its observation [105], through oculomotor control [106, 107] to communication [108, 109]. This application of PC rests on the simple observation that perception and action are in the service of minimizing variational free energy, and action can be cast as minimizing (proprioceptive) prediction errors by changing (sensory) data. Machine learning implementations of PC are now starting to move beyond classification and generative modeling of static images to address dynamic control and reinforcement learning [110, 111, 112]. Active NGC (ANGC) [112] and active PC (ActPC) [113] formulate active inference in terms of predictive processing, much in contrast to machine learning implementations with deep neural networks trained via backprop. ANGC focused on a simple modular agent design, implementing a policy circuit and a generative transition dynamics that learned from sparse reward feedback by using the transition dynamics to produce a dynamically normalized epistemic signal that promoted intelligent exploration. ActPC builds on ANGC by implementing a policy circuit, an actor circuit, a generative transition dynamics circuit, and a prior preference circuit, further exploring how the temporal processing ability of each NGC component could be improved by integrating a simple form of working memory into the neural dynamics. ActPC and ANGC have been empirically shown to work well on standard reinforcement learning control problems and, more importantly, in the context of (simulated) robotic control [113], of which results for general PC also exist [114, 115, 116].
Parallel (independently developed) and homonym frameworks include active PC (APC) [117] and active PC networks (APCNs) [118], developed to learn part-whole hierarchies from natural images (tackling the difficult part-whole and reference frames learning problems), resulting in an interesting model of active, internally directed perception, i.e, the neural circuitry actively samples the natural image space to extract information for crafting internal representations. Notably, the more general APC model of [117] further offered a potential means of tackling the integrated state-action hierarchy learning problem, conducting a form of hierarchical reinforcement learning. The APC model was applied to a multi-room grid-world and was shown to work well. However, APC/APCNs utilize backprop in portions of their underlying architectures, having to train/adapt locally deep MLPs and recurrent networks (and deep hyper-networks) in order to obtain a desirable performance. ActPC, in contrast, rigidly commits to the principle of locality in terms of both inference and learning.
Beyond control agent design, a cognitive architecture known as the COGnitive Neural GENerative system (CogNGen) [87, 86], built on the ideas inherent to ActPC, has been proposed for tackling reinforcement learning problems. CogNGen instantiates elements of the Common Model of Cognition [119] and combines NGC with hyperdimensional / vector-symbolic models of human memory, implementing a module for the motor cortex, procedural memory, working memory neural buffers, and episodic declarative memory. CogNGen has been shown to perform well on maze-like exploration tasks [86, 120], including those that require maintaining memory within and across episodes.
@@
Cognitive Control / Robotics Reinforcement learning, Robotic control, maze navigation [115, 110, 111, 112, 118, 87] [116, 86, 113, 117, 120]
@@ 5 Plausibility of Predictive Coding in Neuroscience
A common question in machine learning is: at what point can a particular algorithm be considered to be biologically plausible? This arises from the fact that no computer simulation can fully replicate the intricate workings of the brain in every respect, and, hence, there will invariably be certain nuances that render the simulation to be implausible in some way. Furthermore, different research agendas consider different properties for differentiating biologically plausible and non-plausible models. In this section, we discuss the key properties of PC that satisfy this distinction and which do not.
Predictive Coding at Scale.
Broadly speaking, it can be said that PC already occurs at the level of individual brains. This notion is related to the Bayesian brain hypothesis [123, 124, 125, 126, 127, 128], a theory which states that our brain encodes a generative model of the world and perception is nothing more than a form of hypothesis testing [129, 130, 131, 132]. The inversion of this generative model conforms to the free energy principle, which can be read as the brain trying to minimise prediction errors, or variational free energy, given observations of the sensed world [38]. But how does this concept scale when we consider smaller “things”? Do individual brain regions perform PC on their own, and do the individual neurons also act so as to maximize the evidence of their own internal model of neighboring neurons [133]? Answering the above questions could serve as an inspiration to design neural systems and architectures that delegate error corrections across scale. In what follows, we will discuss the biological plausibility of PC at the neuronal level. However, we remark that developing a theory as well as empirical computational models of different scales of PC would prove invaluable.
@@
Our understanding of how PC might be implemented at the neuronal level has certainly changed over the last decade. The initial assumption was that the brain would encode two families of neurons/structures: one in charge of propagating predictions and one in charge of propagating errors [20, 46]. As of today, we have no definitive empirical evidence of the existence of single error neurons, though there is ample evidence for laminar-specific segregation of neuronal populations that may communicate predictions and errors, respectively [134, 135]. In addition, one could read the activity of midbrain dopamine neurons as reporting a certain kind of prediction error [136, 137]. Recent work has shown that error signals could potentially be computed by the local voltage dynamics in the dendrites [138, 139, 140]. For a more detailed description in the context of a possible biological neural implementation, see [141, 140]. Although existing technologies make it difficult to empirically demonstrate PC at the level of a single neuron, we have stronger evidence of PC at the level of neural populations, or brain regions [142]. PC can account for multiple brain phenomena, such as end-stopping and extra-classical receptive fields effects in V1 [20], bistable perception [143], illusory motion [144], repetition-suppression [145], and attentional modulation of cortical processing [146, 147]. It has also been shown that, when listening to speech, the human brain stores acquired information in a hierarchical fashion, where frontoparietal cortices predict the activities of higher level representations [148].

### v3

@@ 3.2 Neural Generative Coding
Figure 5: On the left is an arbitrary neural circuit (non-linearities are omitted for simplicity) that could be constructed under one of the three PC frameworks surveyed. Rao & Ballard PC (PC) only employs generative synapses $\mathbf{W}^{\ell}$, while NGC further employs separate feedback synapses (which may or may not correspond symmetrically to the generative pathway). Note that NGC/BC-DIM notably often employ skip connections ($\mathbf{W}^{7}$).
Despite being developed to model information processing in different brain areas, the above formulation is still biologically implausible, the main implausibility being symmetric weight connections. In what follows, we survey a generalization of PC, to arbitrary wiring patterns, called neural generative coding (NGC). This variation has its roots grounded in cable theory [88, 89] and neuronal compartments [90, 88, 89]. Like PC, NGC employs predict-then-correct learning but incorporates additional neurobiological mechanisms including lateral competition, learned precision-weighting, and membrane potential leakage. The key innovation of NGC is decoupling the forward generative pathway (matrices $\mathbf{W}^{\ell}$) from the feedback pathway (matrices $\mathbf{E}^{\ell}$). This allows flexible connectivity patterns, including skip connections, without requiring that feedback structures mirror the generative model in reverse [24]. Each NGC circuit processes input over $T$ time steps, where $g^{\ell}$ represents a precision-weighting function, $\beta$ controls the update rate, $\gamma$ represents membrane leakage, $f_{D}$ is a dendritic processing function, $\Delta t$ is the time step size, and $\Phi(\mathbf{x}^{\ell}_{t})$ captures lateral interactions. Predictions are computed and neural activities are updated according to:
$\mathbf{\bar{x}}^{\ell}_{t}=g^{\ell}\Big(\mathbf{W}^{\ell+1}\cdot\phi^{\ell+1}(\mathbf{x}^{\ell+1}_{t})\Big),$ (9)
$\mathbf{x}^{\ell}_{t+\Delta t}=\mathbf{x}^{\ell}_{t}+\beta\Big(-\gamma\mathbf{x}^{\ell}_{t}-\mathbf{e}^{\ell}_{t}+(\mathbf{E}^{\ell}\cdot\mathbf{e}^{\ell-1}_{t})\odot f_{D}(\mathbf{x}^{\ell}_{t})+\Phi(\mathbf{x}^{\ell}_{t})\Big).$ (10)
Error signals are calculated as precision-weighted differences between actual and predicted activities, modulated by cross-correlational matrices $\mathbf{\Sigma}^{\ell}$. After settling for $T$ steps, both generative and feedback synapses are updated via multi-factor Hebbian rules with $\gamma_{e}<1$ ensuring feedback synapses evolve more slowly than generative ones:
$\Delta\mathbf{W}^{\ell}=\mathbf{e}^{\ell}_{t}\cdot(\phi^{\ell+1}(\mathbf{x}^{\ell+1}_{t}))^{T},$ (11)
$\Delta\mathbf{E}^{\ell}=\gamma_{e}(\phi^{\ell+1}(\mathbf{x}^{\ell+1}_{t})\cdot(\mathbf{e}^{\ell}_{t})^{T}).$ (12)
For a detailed explanation of all the mechanisms of NGC, we refer to the supplementary material.
@@ 4 Predictive Coding in Machine Learning
Despite the limitations regarding the scale highlighted in the previous chapter, PC has demonstrated remarkable versatility across a large number of machine learning tasks. While early implementations focused primarily on simple classification tasks, recent advances have shown that PC can effectively tackle increasingly sophisticated problems that were previously dominated by backpropagation-based approaches. These developments that we now survey, highlight both the practical potential and current limitations of predictive coding as a comprehensive learning paradigm for modern machine learning applications.
@@
The first application of PC to supervised learning involved training a small PC network to perform image classification on the MNIST dataset, achieving test and train errors comparable to those of an MLP of the same complexity (depth and width) [53]. Since then, similar results have been achieved on convolutional networks trained on datasets of RGB images, such as CIFAR10 and SVHN [57]. In this work, the authors showed that updating the weight parameters alongside the neural activities (i.e., running Equations 5 and 6 in parallel), instead of waiting for the inference phase to converge, improves test accuracies as well as ensures better convergence. These results also extend to structured datasets and graph neural networks, where PC is again able to match the performance of backprop on different benchmarks, with the advantage of learning models that are better calibrated and more robust to adversarial examples [60].
@@
Gaussian assumptions (that underwrite PC) can be limiting in scenarios where we need to model different distributions, such as categorical distributions or mixture models. One such scenario is found in transformer models [91]: the attention mechanism encodes a categorical distribution, computed via the softmax activation. To this end, it is possible to generalize the definition of the energy of a specific layer, originally defined as the squared distance between the predictions $\mathbf{u}^{\ell}$ and the activities $\mathbf{x}^{\ell}$, to the KL divergence between two probability distributions, with predictions and activities serving as sufficient statistics. In detail, the PC model now has the following variational free energy:
@@
where $\mathcal{X}^{\ell}(x)$ is a probability distribution defined for every layer with sufficient statistics $x$. If every distribution is a Gaussian, this is equivalent to the energy defined in Equation 4. This generalization allows predictive-coding-based transformers to perform almost as well as standard transformers with the same model complexity [49]. In computational neuroscience, the use of hybrid generative models (with discrete and continuous state spaces) is fairly established: see, e.g., [92] for a first principles account, which can be regarded as a generalization of PC within the setting of well-defined hierarchical generative models with mixed discrete and continuous states.
@@
Figure 6: Various systems based on NGC-based PC: (a) a temporal NGC model (orange diamonds depict error nodes, grey circles depict state nodes, while arrows represent synaptic projections) for processing temporally-varying sequences of data patterns, e.g., frames from the moving MNIST dataset, (b) a convolutional NGC model (figure adapted from [93]; green grid blocks represent state feature maps, blue grid blocks represent prediction maps, yellow-orange grid blocks represent error maps, and blue parallelograms depict deconvolutional pathways) for extracting distributed representations of natural images, and (c) the CogNGen architecture (figure adapted from [94]), a modular control system that implements several core elements of the Common Model of Cognition [95] with PC and vector-symbolic memory ($\mathbf{r}$ depicts a reward, while $\mathbf{a}$ depicts an action vector).
@@
When processing images, PC has shown promise, particularly when generalizations that utilize convolutional operators are considered to tackle problems in vision such as object recognition and discrimination (using over-complete sparse representations) [96, 97, 98, 99, 100]. Early NGC work used simplifying assumptions like greyscale images and low scene complexity with easily identifiable objects. NGC circuits can process image patches ($P\times P$ pixel grids) to learn low-level patterns such as edges and strokes. NGC was later generalized to natural images through convolutional neural generative coding (Conv-NGC), which integrates standard convolution/deconvolution operations [93]. In Conv-NGC, forward and feedback synapses manipulate feature maps coupled to error units that compute mismatch signals for iterative inference. Notably, Conv-NGC models learn feature representations that implicitly embody image pyramids [101].
@@
PC has achieved a discrete success in temporal sequence modeling. Early work developed hierarchical neural generative modeling for natural video data [102, 99, 103, 104], using top-down information to modulate lower-layer activities for locally invariant representations. This built on PC formulated as a Kalman filter [105, 106] and deep hierarchies of local recurrent networks [107, 108]. One of the earliest NGC formulations addressed time-varying data through the temporal neural coding network (TNCN) [109], later generalized to the parallel temporal neural coding network (P-TNCN) [76]. Both models processed pattern sequences without temporal unfolding, making them local in space and time. They demonstrated zero-shot generalization to unseen sequences, with TNCN adapting to bouncing ball videos of different object quantities and P-TNCN synthesizing frames with varying quantities and qualities without altering synaptic weights. Recently, dynamic PC [110] combined deep learning tools with mixture models, acquiring bio-plausible receptive fields for sequential visual perception tasks.
@@
Another promising aspect of P-TNCNs described above is its ability to conduct continual sequential learning [76], as its generative ability on previously seen sequence modeling tasks did not deteriorate nearly as much as in recurrent networks. Additional effort [75] explored strengthening the memory retention ability of NGC by examining the challenging problem of online cumulative learning, where datasets (or tasks) were presented to the system in the form of a stream and with no indication of when the task was switched. Surprisingly, the NGC-based model was shown to outperform or be competitive with a wide swath of backprop-based approaches that relied on memory-buffers (replay), regularization, and/or auxiliary generative modeling in order to preserve prior knowledge. Stronger performance was seen when another neural circuit drove the lateral recurrent synapses of S-NCN based on competitive learning. This simple task mediating circuit was later improved in follow-up effort [111]. Another later effort developed a stochastic implementation of PC with readable/writable memory (BayesPCN) [67] that was shown to be robust to sample-level forgetting.
Active Inference and Control.
In computational neuroscience, early formulations of active inference were based on equipping PC with reflexes in order to simulate a variety of behaviors; ranging from handwriting and its observation [112], through oculomotor control [113, 114] to communication [115, 116], inspiring works that merge it with kinematic control [117, 118]. Active NGC (ANGC) [119] and active PC (ActPC) [120] formulate active inference through predictive processing, contrasting with backprop-trained networks. ANGC implements modular agent design with policy and transition dynamics circuits learning from sparse rewards, while ActPC adds actor and prior preference circuits with working memory integration. Both work well on standard RL control and simulated robotics [121, 122, 123]. Parallel frameworks include active PC (APC) [124] and active PC networks (APCNs) [125] for learning part-whole hierarchies from natural images and hierarchical reinforcement learning in multi-room grid-worlds. However, these utilize backprop in portions of their architectures. The COGnitive Neural GENerative system (CogNGen) [95, 94, 126] combines NGC with hyperdimensional memory models, implementing motor cortex, procedural memory, working memory buffers, and episodic memory modules, performing well on maze exploration tasks requiring cross-episode memory.
@@
Control / Robotics Reinforcement learning, Robotic control, maze navigation [122, 128, 129, 119, 125, 95] [123, 94, 120, 124, 126]
@@ 5 Plausibility of Predictive Coding in Neuroscience
A common question in machine learning is: at what point can a particular algorithm be considered to be biologically plausible? This arises from the fact that no computer simulation can fully replicate the intricate workings of the brain in every respect, and, hence, there will invariably be certain nuances that render the simulation implausible in some way. Furthermore, different research agendas consider different properties for differentiating biologically plausible and non-plausible models. In this section, we will start by addressing the neuroscientific debate and problems with testing PC and then discuss the key properties of PC that satisfy this distinction and which do not.
Neuroscientific Debate over Predictive Coding.
Despite the fact that PC is a subset of variational Bayesian methodology, there is far less direct experimental support for PC compared to the broader Bayesian brain hypothesis. While there is abundant evidence for the Bayesian brain hypothesis itself [130, 131, 132, 133, 134, 135], direct experimental support for PC across modalities remains more limited [136]. Neuroimaging does support some elements of PC [137, 138, 139], particularly through the framing of mismatch negativity [140, 141]. However, mismatch negativity can also be explained via neural adaptation without invoking hierarchical Bayesian mechanics [136]. Evidence has emerged indicating that PC might be better considered as a cognitive, higher-level explanation rather than a lower-level sensory-based one [136]. Although PC can account for various neurophysiological phenomena such as bistable perception [142], perceptual illusions [143], and properties of V1 neuronal responses [144], fewer studies explicitly test hypotheses inherent to PC. These core hypotheses include expectation-scaled error-signaling neural responses and top-down signaling representing sensory prediction. PC has been criticized as difficult to falsify [145].
The translation of PC from algorithmic specification to biophysical instantiation is often unclear, creating further ambiguity about which implementation is being tested in particular studies. Some PC variants better explain certain neurophysiological data over others [146], making it difficult for neuroscience researchers to test its validity [138]. Some studies suggest that PC is more likely to occur in certain brain regions depending on the sensory modalities being processed [147]. More importantly, there appears to be no evidence of a distinction between predictions and errors at the network level [147, 148]. Nevertheless, despite this unsettled debate surrounding PC and its empirical grounding, advances in neurophysiological methodology offer promise. Higher resolution fMRI, calcium imaging, and other techniques, when paired with anatomical models and simulation, might enable future derivations of finer-grained hypotheses. These advances could also facilitate comparisons across variant PC implementations that share the same underlying theoretical framework, including those with different forms of encoding or representation of prediction errors.
@@
Our understanding of how PC might be implemented at the neuronal level has certainly changed over the last decade. The initial assumption was that the brain would encode two families of neurons/structures: one in charge of propagating predictions and one in charge of propagating errors [20, 53]. As of today, we have no definitive empirical evidence of the existence of single error neurons, though there is ample evidence for laminar-specific segregation of neuronal populations that may communicate predictions and errors, respectively [149, 150]. Recent work has shown that error signals could potentially be computed by the local voltage dynamics in the dendrites [151, 54, 152, 153]. For a more detailed description in the context of a possible biological neural implementation, see [55, 152]. Although existing technologies make it difficult to empirically demonstrate PC at the level of a single neuron, we have stronger evidence of PC at the level of neural populations, or brain regions [138, 136]. PC can account for multiple brain phenomena, such as end-stopping and extra-classical receptive fields effects in V1 [20], bistable perception [154], illusory motion [143], repetition-suppression [155], and attentional modulation of cortical processing [156, 157]. It has also been shown that, when listening to speech, the human brain stores acquired information in a hierarchical fashion, where frontoparietal cortices predict the activities of higher level representations [139].